# Constraint on Neutrino Statistics from Cosmological Data

### v1

@@ 4 Results and Discussions
The grey curves and grey shaded areas indicate the fitting results for CMB and the red ones indicate results for CMB+BAO. The 1D marginalized density of $\kappa_{\nu}$ shows that purely bosonic neutrino can be excluded at 95% CL by CMB. Although more weakly, $\kappa_{\nu}=-1$ can also be excluded when combined with DESI BAO distance measurements. Both of the improvements achieved can be attributed to the more accurate datasets and more stringent constraints on $H_{0}$.
@@
It is important to notice that the correlation between $H_{0}$ and $\kappa_{\nu}$ is enhanced for CMB+BAO. This effect may be understood by the stronger constraints on $\sum m_{\nu}$. For CMB+BAO, the constraint on $\sum m_{\nu}$ is strengthened and a smaller neutrino mass $\sum m_{\nu}\lesssim 0.1$ eV is preferred, as shown in the plot of 1D marginalized density of $\sum m_{\nu}$ in the right column of Fig. 5. This makes both the $k_{\nu}-\sum m_{\nu}$ correlation and the $H_{0}-\sum m_{\nu}$ correlation to a narrower region with smaller $\sum m_{\nu}$, as can be seen by comparing the grey area with the red area in the lower panel of the first column and the second column in Fig. 5. As discussed before, small neutrino mass, say $\sum m_{\nu}\lesssim 0.1$ eV, would have very small effects on the evolution of the universe. So this would break the degeneracy not only between $\sum m_{\nu}$ and $\kappa_{\nu}$ but also between $\sum m_{\nu}$ and $\kappa_{\nu}$. Consequently, the correlation between $\kappa_{\nu}$ and $H_{0}$ is enhanced.
@@
To further investigate the impact of neutrino mass on $\kappa_{\nu}$, we repeat the fitting with $\sum m_{\nu}=0.06\ \mathrm{eV}$ chosen fixed, and then compare this fitting result with the fitting result with variable neutrino mass. The comparison is made in Fig. 6. The results obtained for the CMB dataset are shown in Fig. 6(a) and the results obtained for the CMB+BAO dataset are shown in Fig. 6(b). The red curves and shaded areas denote the fitting results of variable neutrino mass model; the blue ones denote the results of $\sum m_{\nu}=0.06$ eV fixed model.
@@
However, this effect becomes less pronounced after incorporating the BAO data. As seen in the right panel of Fig. 6, the 2D comparison plots for $\kappa-H_{0}$ degeneracies in the two analyses—one with a fixed neutrino mass and the other allowing the mass to vary—are quite similar. The reason is that the DESI BAO measurement favors smaller neutrino mass, as can be seen in the plot of 1D marginalized density of $\sum m_{\nu}$ in Fig. 5. As discussed earlier, low-mass neutrinos become non-relativistic at very late times, and their impact on $\theta_{s}$ is negligible. As a result, whether $\sum m_{\nu}$ slightly varies in a small mass regime or not would not significantly affect $\theta_{s}$. That’s why the degeneracies in the two analyses—one with a fixed neutrino mass and the other allowing the mass to vary—are quite similar.
@@
Table I: Fitting results for the cosmological parameters $\kappa_{\nu}$ and $\sum m_{\nu}$ The second and third column show the mean value of $\kappa_{\nu}$ and the lower bound on $\kappa_{\nu}$ at 95% confidence level. The last two columns provide the mean value of $\sum m_{\nu}$ and its upper bound at 95% confidence level.
$\kappa_{\nu}$ $H_{0}\,[\mathrm{km}\,\mathrm{s}^{-1}\,\mathrm{Mpc}^{-1}]$ $\sum m_{\nu}$[eV] $\chi^{2}$ $\chi^{2}_{\mathrm{CMB}}$ $\chi^{2}_{\mathrm{BAO}}$
CMB+$m_{\nu}$ 0.60 67.92 0.082 2796.13 2796.13 /
CMB+BAO+$m_{\nu}$ 0.93 68.36 0.010 2809.97 2795.44 14.53
CMB+BAO+$H_{0}$+$m_{\nu}$0.17 69.74 0.017 5258.14 2799.23 13.03
CMB+$m_{\nu}=0.06$ eV 0.41 67.87 0.06 2796.14 2796.14 /
CMB+BAO+$m_{\nu}=0.06$ eV 0.98 68.38 0.06 2812.85 2798.80 14.05
Table II: Best-fit parameter values and corresponding $\chi^{2}$ values for different datasets. The columns include the values of $\kappa_{\nu}$, $H_{0}$, the sum of neutrino mass $\sum m_{\nu}$, and the $\chi^{2}$ values for the total dataset, CMB, and BAO data.
The constraints on the neutrino statistical parameter $\kappa_{\nu}$ and neutrino mass $\sum m_{\nu}$ from different datasets are summarized in Table  I. All the fitting results rule out the possibility of purely bosonic neutrinos at 95% confidence level except the one using CMB+BAO+$H_{0}$ dataset, whose fitting quality is poor due to the Hubble tension. Neutrino with mixed statistics remains as a viable option with $-0.5\lesssim\kappa_{\nu}<1$ at $2\sigma$ confidence interval. Comparing the upper bound of $\sum m_{\nu}$ given by CMB+BAO+$m_{\nu}$ fitting to the results in Table 4 of [5], it can be found that introducing mixed statistics for neutrinos causes the DESI data to favor a slightly larger neutrino mass. The upper bound increases from 0.072 eV to 0.078 eV. This helps reduce the potential neutrino mass tension between cosmological observations and neutrino oscillation experiments, but the effect remains minimal.
Table II presents the results of best-fit parameters for each dataset, along with their corresponding $\chi^{2}$ values. All cases with $\kappa_{\nu}>0$ indicate a preference for fermionic neutrinos, especially in the case of CMB+BAO, where, despite a slight increase in $\chi^{2}$ , the preference for fermionic statistics becomes even more evident. The high $\chi^{2}$ value in the fourth row indicates that, in the mixed statistics neutrino model, the Hubble tension issue remains unresolved.
@@ 5 Conclusions
We fit the variable-neutrino-statistics model with CMB, CMB+BAO and CMB+BAO+$H_{0}$ datasets respectively. We find that purely bosonic neutrinos have been completely excluded by the CMB or CMB+BAO datasets. This is our first key conclusion.
Based on the mean values and the best fit results of $\kappa_{\nu}$, we conclude that the cosmological data strongly favor neutrinos following Fermi-Dirac statistics, while neutrinos with mixed statistics remain a possibility. This forms our second key conclusion.
We further fit the variable-neutrino-statistics model with $\sum m_{\nu}=0.06$ eV fixed using CMB and CMB+BAO datasets respectively. By comparing the fitting results with the ones from variable neutrino mass model, we find that the correlation between $\kappa_{\nu}$ and $H_{0}$ is enhanced once $\sum m_{\nu}$ is fixed in the case of CMB. This effect is less pronounced in the case of CMB+BAO.

### v2

@@ 4 Results and Discussions
The grey curves and grey shaded areas indicate the fitting results for CMB and the red ones indicate results for CMB+BAO. The 1D marginalized density of $\kappa_{\nu}$ shows that purely bosonic neutrino is strongly disfavoured by CMB. Although more weakly, including DESI BAO distance measurements also disfavors $\kappa_{\nu}=-1$. The dashed lines show that purely bosonic neutrino can be excluded at 95% CL by both CMB and CMB+BAO. As shown in the upper panel of the second column in Fig. 5, when BAO distance measurements are included, the fitting would prefer slightly higher $H_{0}$s. The degeneracy between $H_{0}$ and $\kappa_{\nu}$ makes relatively smaller $\kappa_{\nu}$ preferred. Therefore, the constraint from CMB+BAO is weaker than the one from CMB. Both of the improvements achieved here with respect to results in [8] can be attributed to the more accurate datasets and more stringent constraints on $H_{0}$.
@@
It is important to notice that the correlation between $H_{0}$ and $\kappa_{\nu}$ is enhanced for CMB+BAO. Here, ”enhanced correlation” refers to a narrower posterior distribution along the degeneracy direction, reflecting tighter constraints. This effect may be understood by the stronger constraints on $\sum m_{\nu}$. For CMB+BAO, the constraint on $\sum m_{\nu}$ is strengthened and a smaller neutrino mass $\sum m_{\nu}\lesssim 0.1$ eV is preferred, as shown in the plot of 1D marginalized density of $\sum m_{\nu}$ in the right column of Fig. 5. This makes both the $k_{\nu}-\sum m_{\nu}$ correlation and the $H_{0}-\sum m_{\nu}$ correlation to a narrower region with smaller $\sum m_{\nu}$, as can be seen by comparing the grey area with the red area in the lower panel of the first column and the second column in Fig. 5. As discussed before, small neutrino mass, say $\sum m_{\nu}\lesssim 0.1$ eV, would have very small effects on the evolution of the universe. So this would break the degeneracy not only between $\sum m_{\nu}$ and $\kappa_{\nu}$ but also between $\sum m_{\nu}$ and $H_{0}$. Consequently, the correlation between $\kappa_{\nu}$ and $H_{0}$ is enhanced.
@@
To further investigate the impact of neutrino mass on $\kappa_{\nu}$, we perform a fit with $\sum m_{\nu}=0.06\ \mathrm{eV}$ fixed, and compare it to the case with variable neutrino mass. The comparison is made in Fig. 6. The results obtained for the CMB dataset are shown in Fig. 6(a) and the results obtained for the CMB+BAO dataset are shown in Fig. 6(b). The red curves and shaded areas denote the fitting results of variable neutrino mass model; the blue ones denote the results of $\sum m_{\nu}=0.06$ eV fixed model.
@@
However, this effect becomes less pronounced after incorporating the BAO data. As seen in the right panel of Fig. 6, the 2D comparison plots for $\kappa-H_{0}$ degeneracies in the two analyses—one with a fixed neutrino mass and the other allowing the mass to vary—are quite similar. The reason is that the DESI BAO measurement favors smaller neutrino mass, as can be seen in the plot of 1D marginalized density of $\sum m_{\nu}$ in Fig. 5. As discussed earlier, low-mass neutrinos become non-relativistic at very late times, and their impact on $\theta_{s}$ is negligible. As a result, whether $\sum m_{\nu}$ slightly varies in a small mass regime or not would not significantly affect $\theta_{s}$. That is why the degeneracies in the two analyses—one with a fixed neutrino mass and the other allowing the mass to vary—are quite similar.
@@
Table II: Fitting results for the cosmological parameters $\kappa_{\nu}$ and $\sum m_{\nu}$ The second and third column show the mean value of $\kappa_{\nu}$ and the lower bound on $\kappa_{\nu}$ at 95% confidence level. The last two columns provide the mean value of $\sum m_{\nu}$ and its upper bound at 95% confidence level.
$\kappa_{\nu}$ $H_{0}$ $\sum m_{\nu}$[eV] $\chi^{2}$ $\chi^{2}_{\mathrm{CMB}}$ $\chi^{2}_{\mathrm{BAO}}$ $\chi^{2}_{\mathrm{SNe}}$
CMB+$m_{\nu}$ 0.60 67.92 0.082 2796.13 2796.13 / /
CMB+BAO+$m_{\nu}$ 0.93 68.36 0.010 2809.97 2795.44 14.53 /
CMB+BAO+SNe+$m_{\nu}$0.04 69.40 0.034 3852.08 2799.22 13.50 1039.35
CMB+$m_{\nu}=0.06$ eV 0.41 67.87 0.06 2796.14 2796.14 / /
CMB+BAO+$m_{\nu}=0.06$ eV 0.98 68.38 0.06 2812.85 2798.80 14.05 /
Table III: Best-fit parameter values and corresponding $\chi^{2}$ values for different datasets. The columns include the values of $\kappa_{\nu}$, $H_{0}$, the sum of neutrino mass $\sum m_{\nu}$, and the $\chi^{2}$ values for the total dataset, CMB, and BAO data. The reported $\chi^{2}$ values are obtained from the MCMC chains. The Hubble constant $H_{0}$ is expressed in units of km s-1 Mpc-1.
The constraints on the neutrino statistical parameter $\kappa_{\nu}$ and neutrino mass $\sum m_{\nu}$ from different datasets are summarized in Table  II. All the fitting results rule out the possibility of purely bosonic neutrinos at 95% confidence level except the one using CMB+BAO+SNe dataset, whose fitting quality may be poor due to the Hubble tension. Neutrino with mixed statistics remains as a viable option with $-0.5\lesssim\kappa_{\nu}<1$ at $2\sigma$ confidence interval. Comparing the upper bound of $\sum m_{\nu}$ given by CMB+BAO+$m_{\nu}$ fitting to the results in Table 4 of [5], it can be found that introducing mixed statistics for neutrinos causes the DESI data to favor a slightly larger neutrino mass. The upper bound increases from 0.072 eV to 0.078 eV. This helps reduce the potential neutrino mass tension between cosmological observations and neutrino oscillation experiments, but the effect remains minimal.
Table III presents the results of best-fit parameters for each dataset, along with their corresponding $\chi^{2}$ values. According to the table, the results from CMB+BAO+$m_{\nu}$ and CMB+BAO+ $m_{\nu}=0.06$ eV combinations indicate that fermionic statistics is preferred by CMB+BAO. In contrast, the results of CMB+$m_{\nu}$ and CMB+$m_{\nu}=0.06$ eV suggest a preference for a mixed distribution by CMB. The results of CMB+BAO+SNe+$m_{\nu}$ suggest a stronger preference for the mixed statistics neutrino model. However, due to the presence of the Hubble tension in this dataset combination, this result should be viewed with caution and considered for reference only.
@@ 5 Conclusions
We fit the variable-neutrino-statistics model with CMB, CMB+BAO and CMB+BAO+SNe datasets respectively. We find that purely bosonic neutrinos have been excluded by the CMB or CMB+BAO datasets at $2\sigma$. This is our first key conclusion.
Based on the mean values and the best fit results of $\kappa_{\nu}$, we conclude that CMB+BAO dataset favors neutrinos following Fermi-Dirac statistics, while neutrinos with mixed statistics remain a possibility especially when only CMB data are included in the fit. This forms our second conclusion.
@@
Note added
After the submission of this article, the DESI Collaboration made their second data release (DESI DR2) [38]. We have combined the DESI DR2 data with CMB data and repeated the fitting CMB+BAO defined in Section 3 of the main text. We find no significant differences compared to the previous results, as can be seen in Fig. 7 and Fig. 8. Our first conclusion remains valid, while the second one is weakened slightly because the best fit value of $\kappa_{\nu}$ moves to a slightly smaller value, as seen in the third line in Table IV.
[…]
Figure 8: The cosmological constraints for variable-neutrino-statistics model. This figure is identical to Fig. 6 in the main text, except that the BAO data have been replaced with DESI DR2.
$\kappa_{\nu}$ $\sum m_{\nu}$ [eV] $H_{0}$ [km $\mathrm{s}^{-1}$ $\mathrm{Mpc}^{-1}$] $\chi^{2}$ $\chi^{2}_{\mathrm{CMB}}$ $\chi^{2}_{\mathrm{BAO}}$
CMB+DESI DR1 0.93 0.010 68.36 2809.97 2795.44 14.53
CMB+DESI DR2 0.76 0.018 68.64 2807.00 2795.34 11.66
CMB+DESI DR1+$m_{\nu}=0.06$ eV 0.98 / 68.38 2812.85 2798.80 14.05
CMB+DESI DR2+$m_{\nu}=0.06$ eV 0.45 / 68.94 2812.45 2801.31 11.14
Table IV: Comparison of the best-fit values of $\kappa_{\nu}$, $\sum m_{\nu}$ and $H_{0}$.
@@ Acknowledgements
The authors are grateful to the anonymous referee for the careful reading of the manuscript and many insightful suggestions.